A pension fund manager is considering three mutual funds. The first is a stock fund, the second is a long-term government and corporate bond fund, and the third is a T-bill money market fund that yields a sure rate of 4.0%. The probability distributions of the risky funds are:

Expected Return Standard Deviation
Stock fund (S) 10% 32%
Bond fund (B) 7% 24%


The correlation between the fund returns is 0.1250.

What is the Sharpe ratio of the best feasible CAL?
Answer #1

Expected return stock = E(R1 )

Expected return bond = E(R2 )

Standard deviation sock= σ1

Standard deviation bond= σ2

weight of stock = w1

weight of bond = w2

The expected return of the risky portfolio= w1E(R1) + w2E(R2)

the covariance of risky portfolio= w21σ21 + w22σ22 + 2ρ(R1, R2) w1w2σ1σ2, using ρ(R1, R2), the correlation of R1 and R2

hence, we get the following risk weights and retuns for the risky portfolio as shown in Table 1

Now we compute the complete portfolio which

The expected return of a complete portfolio is given as:

E(Rc) = wpE(Rp) + (1 − wp)Rf (see table 2 for computation of risk free returns based on weights )

And the variance and standard deviation of the complete portfolio return is given as:

Var(Rc) = w2pVar(Rp), σ(Rc) = wpσ(Rp) (see table 2 for computation of risk free returns based on weights )

The line E(Rc) = Rf + Spσ(Rc) is the capital allocation line (CAL). The slope of the line, Sp, is called the Sharpe ratio

hence, by replacing the values in the formula we get

6.25%=4%+Sp*2.24%

i.e Sp= 1.00446429

Table 1

weight of stocks Weight of bonds Riski portfolio standard deviation Expected return
0.1 0.9 4.94% 7.30%
0.2 0.8 4.40% 7.60%
0.3 0.7 4.15% 7.90%
0.4 0.6 4.17% 8.20%
0.5 0.5 4.48% 8.50%
0.6 0.4 5.07% 8.80%
0.7 0.3 5.94% 9.10%
0.8 0.2 7.09% 9.40%
0.9 0.1 8.52% 9.70%
1 0 10.24% 10.00%

Table 2

riski portfolio weights Risk free assets weights 0.049408 0.044032 0.041472 0.041728 0.0448 0.050688 0.059392 0.070912 0.085248 0.1024
0.1 0.9 4.33% 4.33% 4.33% 4.33% 4.33% 4.33% 4.33% 4.33% 4.33% 4.33%
0.2 0.8 4.72% 4.72% 4.72% 4.72% 4.72% 4.72% 4.72% 4.72% 4.72% 4.72%
0.3 0.7 5.17% 5.17% 5.17% 5.17% 5.17% 5.17% 5.17% 5.17% 5.17% 5.17%
0.4 0.6 5.68% 5.68% 5.68% 5.68% 5.68% 5.68% 5.68% 5.68% 5.68% 5.68%
0.5 0.5 6.25% 6.25% 6.25% 6.25% 6.25% 6.25% 6.25% 6.25% 6.25% 6.25%
0.6 0.4 6.88% 6.88% 6.88% 6.88% 6.88% 6.88% 6.88% 6.88% 6.88% 6.88%
0.7 0.3 7.57% 7.57% 7.57% 7.57% 7.57% 7.57% 7.57% 7.57% 7.57% 7.57%
0.8 0.2 8.32% 8.32% 8.32% 8.32% 8.32% 8.32% 8.32% 8.32% 8.32% 8.32%
0.9 0.1 9.13% 9.13% 9.13% 9.13% 9.13% 9.13% 9.13% 9.13% 9.13% 9.13%
1 0 10.00% 10.00% 10.00% 10.00% 10.00% 10.00% 10.00% 10.00% 10.00% 10.00%

Table 3

riski portfolio weights Risk free assets weights Riski portfolio standard deviation Risk of risk free portfolio
0.1 0.9 4.94% 0.49%
0.2 0.8 4.40% 0.88%
0.3 0.7 4.15% 1.24%
0.4 0.6 4.17% 1.67%
0.5 0.5 4.48% 2.24%
0.6 0.4 5.07% 3.04%
0.7 0.3 5.94% 4.16%
0.8 0.2 0.070912 5.67%
0.9 0.1 0.085248 7.67%
1 0 0.1024 10.24%

Table 1